OpenStudy (ladiesman217):

I GIVE MEDALS AND FANS help me please Part 1: Using complete sentences, explain how to find the equation of the line, in standard form and slope–intercept form, passing through (3, 6) and (–2, –4). Part 2: Compare the benefits of writing an equation in standard form to the benefits of writing an equation in slope–intercept form.

OpenStudy (anonymous):

standard form is Ax +By = C; slope intercept form is y = mx + b i would find the slope first m = (-4-6)/(-2-3) = -10/-5 = 2 slope intercept form would be y = 2x + b (-4 - b)/(-2-0) = 2; solve for b -4-b= -4, b=0 y = 2x Slope intercept 2x - y = 0 Standard form

OpenStudy (anonymous):

part 2: im not entirely sure what would be considered benefits except in slope intercept you can see where x =0 intercepts the y axis and what the slope of the line is. and in standard you have variables all on one side and constants on the other

OpenStudy (ladiesman217):

ok thanks